The company has two departments – Cutting and Sewing. The company uses the multiple production department factory overhead rate method using direct labor hours. The budgeted factory overhead is $1,120,000 for the Cutting Department and $660,000 for the Sewing Department. The company plans to make 100,000 shirts and 60,000 pairs of pants. It takes 1 hour to cut and 1 hour to sew a shirt. It takes 1 hour to cut and 2 hours to sew a pair of pants. How many direct labor hours are being worked in each department? What is the factory overhead rate per direct labor hour for each department?

Direct labor hours Cutting Dept?

Direct Labor hours Sewing Dept?

Factory Overhead Rate for the Cutting Dept?

Factory Overhead Rate for the Sewing Dept?

How much FOH is allocated to a single shirt?

How much FOH is allocated to a single pair of pants?
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
The company
Answer 1 Shirt Pants Total Note
Units produced      100,000.00      60,000.00 A
Time taken to cut 1 unit                  1.00                1.00 B
Direct labor hours in Cutting Dept.      100,000.00     60,000.00 160,000.00 C=A*B
Answer 2 Shirt Pants Total Note
Units produced      100,000.00      60,000.00 See A
Time taken to sew 1 unit                  1.00                2.00 D
Direct labor hours in Sewing Dept.      100,000.00 120,000.00 220,000.00 E=A*D
Answer 3 and 4 Cutting Sewing Note
Direct labor hours      160,000.00 220,000.00 F=C,E
Budgeted factory overhead 1,120,000.00 660,000.00 G
Factory Overhead Rate                  7.00                3.00 H=G/F
Answer 5 Cutting Sewing Total Note
Time taken for 1 unit                  1.00                1.00 I=B,D
Departmental predetermined OH Rate                  7.00                3.00 See H
FOH allocated to a single shirt                  7.00                3.00             10.00 J=I*H
Answer 6 Cutting Sewing Total Note
Machine hours used                  1.00                2.00 K=B,D
Departmental predetermined OH Rate                  7.00                3.00 See H
FOH allocated to a single pair of pants                  7.00                6.00             13.00 L=K*H